OpenStudy (anonymous):

How many zeros will be in the standard form of the exponential expression ten to the ninety-ninth power? ten zeros ninety-eight zeros ninety-nine zeros one hundred zeros

OpenStudy (anonymous):

So, do you have an idea of how to start this?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Alright then. we'll do a similar kind of problem that perhaps isn't quite so large. 10^2 = 100. The exponent will dictate how many zeros you have after the 1. Make sense?

OpenStudy (anonymous):

Because 10 multiplied by 10 is 100.

OpenStudy (anonymous):

Or 10^3, will have 3 zeros for 1000.

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

If the exponent in your problem is 99, how many zeros do you think your problem will have?

OpenStudy (austinl):

Think of it like this, \(\large{10^{99}\Rightarrow (10)_1(10)_2(10)_3\cdot\cdot\cdot(10)_{99}}\) When you multiply anything by 10, you simply add on a zero to the end. Knowing this, what do you think is the answer?

OpenStudy (anonymous):

so its 99

OpenStudy (anonymous):

Right. :)

OpenStudy (anonymous):

thanks guys love u i will give u a ward award

OpenStudy (anonymous):

can celecity can u answer my other question plzz

OpenStudy (anonymous):

Post the link for me and I'll take a look at it. :) I'll see what I can do

OpenStudy (anonymous):

Suppose your friend starts a rumor by telling you some juicy gossip Sunday night. On Monday, one of you shares the gossip with 2 more friends. The next day, they each share the rumor with 2 of their friends, so 4 more people hear the rumor on Tuesday. On Wednesday, those people each share the rumor with 2 of their friends, so now 8 more people have heard the rumor. If the rumor continues to spread this way, how many people will hear the rumor on Saturday? sixteen thirty-two sixty-four one hundred twenty-eight

OpenStudy (anonymous):

Oh okay. So let's make a table so we can see these numbers better. Sound cool?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

 The numbers are the amount of people know the rumor on each day. What do you notice about these numbers?